hello,
once two equations are in slope-intercept form, are they only solvable graphically or can you solve them algebraically or otherwise?
for instance, Y=1/6x + 3 and Y=2/3 + 0. Their answer is 6,4. Please show me the steps to this answer.
thank yous.

Answer by Alan3354(69443)

once two equations are in slope-intercept form, are they only solvable graphically or can you solve them algebraically or otherwise?
for instance, Y=1/6x + 3 and Y=2/3 + 0. Their answer is 6,4. Please show me the steps to this answer.
-----------
Your 2nd eqn has no x term.
If you meant:
Y = (1/6)x + 3 and Y = (2/3)x + 0
----
Since they both = y,
(1/6)x + 3 = (2/3)x
Multiply by 6
x + 18 = 4x
3x = 18
x = 6
y = 4
--> (6,4)